OpenStudy (pinksapphire):

How do I factor this out and what are the zeros? y=12x^2-4x+6x-2

OpenStudy (anonymous):

,factor out 4 from first two terms 12x^2 - 4x & factor out 2 from 6x-2 What do you get ?

OpenStudy (pinksapphire):

Is this right? 4x(3x-1)+2(3x-1)

OpenStudy (anonymous):

Sorry factor out 4x*

OpenStudy (anonymous):

Yes correct !!

OpenStudy (anonymous):

Now factor out common term again

OpenStudy (pinksapphire):

How?

OpenStudy (anonymous):

See we can factor out (3x - 1) .. right ?

OpenStudy (pinksapphire):

Yes

OpenStudy (anonymous):

So what do you get ?

OpenStudy (pinksapphire):

For the zeros?

OpenStudy (anonymous):

After factor out (3x-1)

OpenStudy (pinksapphire):

-3?

OpenStudy (anonymous):

no please check it after factor out we have (3x-1)*(4x+2) right ?

OpenStudy (pinksapphire):

Can you show me how? Sorry, I started learning this in Algebra today so...

OpenStudy (pinksapphire):

Now, I got it...I thought you had to do something else first.

OpenStudy (pinksapphire):

How do I get the zeros though?

OpenStudy (anonymous):

Okay we have 4x*(3x - 1) + 2*(3x - 1) then we factor out (3x - 1) from this So we get (3x-1)(4x* 1 + 2* 1) that is (3x-1)*(4x+2)

OpenStudy (anonymous):

to get zeros just equate this factor with zero that is (3x - 1) * (4x - 2) = 0

OpenStudy (anonymous):

Now use zero property ...

OpenStudy (pinksapphire):

Can you show me how?

OpenStudy (anonymous):

Zero property if (a) * (b) = 0 then (a)= 0 or (b) = 0

OpenStudy (anonymous):

And then solve for x

OpenStudy (anonymous):

Can you apply zero property ?

OpenStudy (pinksapphire):

No :(

OpenStudy (anonymous):

see we have (3x-1)*(4x +2) =0 so using that property we can write it as (3x-1) = 0 or (4x+2) = 0 Okay ?

OpenStudy (pinksapphire):

So it will be x=1/3 and 1/2

OpenStudy (anonymous):

Please check sign of 1/2 ..

OpenStudy (pinksapphire):

-1/2

OpenStudy (anonymous):

Correct !!

OpenStudy (pinksapphire):

Thanks!

OpenStudy (anonymous):

You are most welcome :)